Please help with the below.

Thank you!

Please use R Console:

I have to create a command in R console so that below all has 3 bins.

Below is the data:

107 194 146 40 8 4 1

Please show the command so that all have minimum 3 bins.

The command should be something like the below:

newbreaks=c(breaks[1], breaks[i:j],breaks[n])
Answer #1


> rm(list=ls())
> (breaks=c(107,194,146,40,8,4,1))
[1] 107 194 146 40 8 4 1
> (n=length(breaks))
[1] 7
> (i=n-(n-2))
[1] 2
> (j=n-1)
[1] 6
> (newbreaks=c(breaks[1],breaks[i:j],breaks[n]))
[1] 107 194 146 40 8 4 1
>